Compute the weighted average cost of capital given: The before tax cost of debt is 10.5% The cost of preferred stock is 11.75% The cost of common stock equity is 13% The tax rate is 40% The sources of capital are: Long term debt 40% Preferred stock 15% Common stock equity ????? Use the following format for your answer: 7.50%
Answer #1

Weight of common stock equity= 1 - (0.40 + 0.15)

  = 1 - 0.55

= 0.45

The weighted average cost of capital is calculated using the below formula:

WACC=Wd*Kd(1-t)+Wps*Kps+We*Ke

where:

Wd= Percentage of debt in the capital structure.

Kd= The before tax cost of debt

Wps= Percentage of preferred stock in the capital structure

Kps=Cost of preferred stock

We=Percentage of equity in the capital structure

Ke= The cost of common equity.

T= Tax rate

WACC= 0.40*10.5%*(1 - 0.40) + 0.15*11.75% + 0.45*13%

= 0.40*6.30% + 0.15*11.75% + 0.45*13%

= 2.52% + 1.7625% + 5.85%

= 10.1325% 10.13%.

In case of any query, kindly comment on the solution.